Express the complex number 6i in polar form, z=r⁢ei⁢θ. Use a value for θ with 0≤θ<2π.

r = ?
Answer #1

\mathrm{Let, \;z= 6i}
\mathrm{\;z= 6(0+i)}
\mathrm{\;z= 6(\cos\frac{\pi}{2}+i\sin\frac{\pi}{2})}
\mathrm{\;z= 6e^{i\frac{\pi}{2}}} \;\;\;\;\;\;\;\;::\mathrm{ {\color{Blue} e^{i\theta} = (\cos\theta+i\sin\theta)}}
Please comment if you have any doubt in the solution, I will be happy to help you!
Happy Learning!!!!!!!